OpenStudy (anonymous):

Which set of points are on the line y = 2x-1

OpenStudy (lucaz):

do you have some options?

OpenStudy (lucaz):

okay, all you have to do is substitte the x value of each point (x,y) and see if you get the corresponding y value, for example the first given point (0,-1) y=2(x)-1 y=2(0)-1 y=0-1 y=-1 the point (0,-1) lies on this line try the other points

OpenStudy (lucaz):

if you do the subtitution process you will find the right set